OpenStudy (anonymous):

Help asap plz ill give a gold medal. the label on cars antifreeze claims to protect the car between -30celsius and 130celsius. to convert Celsius temperature Fahrenheit temperature, the formula is, c=5/9(F-32). Write an solve and inequality to determine the Fahrenheit temperature range at which antifreeze protects the car.

OpenStudy (anonymous):

you have C and you want F use \[F=\frac{9}{5}C+32\] to find it

OpenStudy (anonymous):

now you need to replace \(C\) by \(-30\) and get \[F=\frac{9}{5}\times (-30)+32\] \[F=-54+32=-22\]

OpenStudy (anonymous):

next step is to replace \(C\) by \(130\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

cant figure it out can you keep going

OpenStudy (anonymous):

can you help with this one mertsj

OpenStudy (anonymous):

first step is to write \[F=\frac{9}{5}\times 130+32\]

OpenStudy (anonymous):

then you have to compute \[\frac{9}{5}\times 130\] how many times does \(5\) go in to \(130\) i.e. what is \(130\div 5\)?

OpenStudy (anonymous):

um 26

OpenStudy (anonymous):

ok good now what is \(9\times 26\)?

OpenStudy (anonymous):

234

OpenStudy (anonymous):

right, so \(\frac{9}{5}\times 130=234\) and finally what is \(234+32\)?

OpenStudy (anonymous):

266

OpenStudy (anonymous):

so it is therefor the answer you are looking for is it protects cars from \(-22\) to \(226\) or you could even write \[-22\leq x\leq 226\] if you like